# Kernelized Concept Erasure

### v5

@@ A.3 Experimental setting
The kernelized minimax game.
@@
We approximate the kernel space using $L=1024$ Nyström landmarks. We run the adversarial game Eq. 10 for each of the kernel mappings we consider, by performing alternate minimization and maximization over ${\boldsymbol{\theta}}$ and $P$, respectively. As our optimization procedure, we use stochastic gradient descent with a learning rate of 0.08 and minibatches of size 256. We run for 35,000 batches, and choose the projection matrix $P$ which leads to the biggest decrease in the linear classification accuracy on the evaluation set. In all cases, we identify a matrix which decreases classification accuracy to near-random accuracy. All training is done on a single NVIDIA GeForce GTX 1080 Ti GPU.
@@
We train an MLP with 2 hidden layers of sizes 512 and 300 to map the original inputs $\mathbf{x}_{n}$ to inputs which, after being mapped to kernel space, are close to the neutralized features. We use dropout of 0.1, ReLU activation and layer normalization after each hidden layer. We use a skip connection between the input and the output layer, i.e., we set the final output of the MLP to be the sum of its inputs and outputs. [^13] This is useful because we assume the gender information is not very salient in the original representations, since these encode much more information. With the skip connection, the MLP only needs to learn what to remove, rather than learning to generate plausible representations. We train for 15,000 batches of size 128 and choose the model that yields the lowest loss on the evaluation set.
@@ A.3.1 Pipeline Evaluation
In this appendix, we include sanity check experiments that aim to assess whether the minimax game Eq. 10 effectively removes linearly-present concepts from the non-linear kernel features, and whether the training of the pre-image network succeeds.
@@
Do we effectively neutralize the concept in the approximate kernel space? For each kernel, we solve the game Eq. 10 and use the final projection matrix $P$ to create neutralized features. We neutralize the features in RKHS by mapping $\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})\mapsto P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})$, and train a linear classifier to recover the gender labels from the neutralized representations. We get a classification accuracy of 50.59$\pm$0.04, very close to majority accuracy of 50.58. This suggests that the process is effective in protecting against the kernel which was applied in training (training a linear classifier on the kernel-transformed representations is equivalent to training a kernel classifier on the original representations). Notice, however, that we cannot test other non-linear kernel classifiers on these representations in a similar way: If the approximate kernel mapping $\widetilde{\boldsymbol{\Phi}}(\cdot)$ corresponds, for example, to a polynomial kernel, we cannot measure the success of an RBF kernel in recovering the bias information after the intervention without performing the pre-image mapping.
@@
Our neutralization algorithm relies on calculating the pre-image of the kernel features after the intervention. To evaluate the quality of the pre-image mapping, we measure the relative reconstruction error $\left|\left|\frac{{P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})-\boldsymbol{\widetilde{\Phi}}(f(\mathbf{x}_{n}))}}{P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})}\right|\right|^{2}_{2}$ over all points $\mathbf{x}_{n}$ on the evaluation set. When averaged over all 4 seeds and the different kernels we experimented with, we get a reconstruction error of 1.81$\pm$1.61(range 0.45-7.94).

### v6

@@ A.3 Experimental setting
The kernelized maximin game.
@@
We approximate the kernel space using $L=1024$ Nyström landmarks. We run the adversarial game Eq. 10 for each of the kernel mappings we consider, by performing alternate minimization and maximization over ${\boldsymbol{\theta}}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$, respectively. As our optimization procedure, we use stochastic gradient descent with a learning rate of 0.08 and minibatches of size 256. We run for 35,000 batches, and choose the projection matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ which leads to the biggest decrease in the linear classification accuracy on the evaluation set. In all cases, we identify a matrix which decreases classification accuracy to near-random accuracy. All training is done on a single NVIDIA GeForce GTX 1080 Ti GPU.
@@
We train an MLP with 2 hidden layers of sizes 512 and 300 to map the original inputs $\boldsymbol{x}_{n}$ to inputs which, after being mapped to kernel space, are close to the neutralized features. We use dropout of 0.1, ReLU activation and layer normalization after each hidden layer. We use a skip connection between the input and the output layer, i.e., we set the final output of the MLP to be the sum of its inputs and outputs. [^13] This is useful because we assume the gender information is not very salient in the original representations, since these encode much more information. With the skip connection, the MLP only needs to learn what to remove, rather than learning to generate plausible representations. We train for 15,000 batches of size 128 and choose the model that yields the lowest loss on the evaluation set.
@@ A.3.1 Pipeline Evaluation
In this appendix, we include sanity check experiments that aim to assess whether the maximin game Eq. 10 effectively removes linearly-present concepts from the non-linear kernel features, and whether the training of the pre-image network succeeds.
@@
Do we effectively neutralize the concept in the approximate kernel space? For each kernel, we solve the game Eq. 10 and use the final projection matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ to create neutralized features. We neutralize the features in RKHS by mapping $\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})\mapsto{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})$, and train a linear classifier to recover the gender labels from the neutralized representations. We get a classification accuracy of 50.59$\pm$0.04, very close to majority accuracy of 50.58. This suggests that the process is effective in protecting against the kernel which was applied in training (training a linear classifier on the kernel-transformed representations is equivalent to training a kernel classifier on the original representations). Notice, however, that we cannot test other non-linear kernel classifiers on these representations in a similar way: If the approximate kernel mapping $\widetilde{\boldsymbol{\Phi}}(\cdot)$ corresponds, for example, to a polynomial kernel, we cannot measure the success of an RBF kernel in recovering the bias information after the intervention without performing the pre-image mapping.
@@
Our neutralization algorithm relies on calculating the pre-image of the kernel features after the intervention. To evaluate the quality of the pre-image mapping, we measure the relative reconstruction error $\left|\left|\frac{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})-\boldsymbol{\widetilde{\Phi}}(f(\boldsymbol{x}_{n}))}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})}\right|\right|^{2}_{2}$ over all points $\boldsymbol{x}_{n}$ on the evaluation set. When averaged over all 4 seeds and the different kernels we experimented with, we get a reconstruction error of 1.81$\pm$1.61(range 0.45-7.94).